Determine whether the integral is convergent or divergent. 0 34 e1/x x3 dx −1 convergent divergent If it is convergent, evaluate it. (If the quantity diverges, enter DIVERGES.)

6 + -1 points SCalcET8 7.8.039. 013 Submissions Used Determine whether the integral is convergent or divergent. 1034 1/x convergent divergent If it is convergent, evaluate it. (If the quantity diverges, enter DIVERGES.)
Answer #1

Consider the integral, Put-t Then, dx = dt As x--1, then t->1 AS x → 0 , then t→oo Therefore, -341-e-x-tx(dt) t-l To find theLet u = t, v = e-t . Then, に! に! -34-teedt t=1 34-te-e t=1 -34 te-t + e-t t-1 34 limte limelimte +lime 11 (1 1 34m- 34(1.0)-